OpenStudy (ledah):

identify the x and y intercepts for the equation 3x + 4y = 12. (4, 0) and (0, -3) (3, 0) and (0, -4) (4, 0) and (0, 3) (-3, 0) and (0, 4) (3, 0) and (0, 4)

OpenStudy (anonymous):

x-int: (4,0) y-int:(0,3) just replace y with 0 to get x-int and then replace x with zero to get y-int

OpenStudy (ledah):

does it have to equal to 12?

OpenStudy (anonymous):

to get x-int replace y with zero then solve for x to get the y-int replace x with zero and solvr for y

OpenStudy (anonymous):

to get x-intercet given \[3x+4y=12\] \[3(0)+4y=12\] \[4y=12\] \[y=3\]

OpenStudy (ledah):

oh. I see now. Thanks.

OpenStudy (anonymous):

to get y-intercept given \[3x+4y=12\] \[3x+4(0)=12\] \[3x=12\] \[x=4\]

OpenStudy (anonymous):

welcome